OpenStudy (anonymous):

Solve and check for extraneous solutions. (5x+20)/(x^2+3x) - (x-2)/(x+3) = 6/x

OpenStudy (fibonaccichick666):

that is fun...

OpenStudy (anonymous):

??????

OpenStudy (fibonaccichick666):

ok so first. what do you need to add or subtract fractions?

OpenStudy (anonymous):

a common denominator

OpenStudy (fibonaccichick666):

ok so (this is only one way of going about this) what is the LCM of (x+3) and \((x^2+3x)\)

OpenStudy (anonymous):

x+3

OpenStudy (fibonaccichick666):

ok so now, can you tell me how you got that?

OpenStudy (anonymous):

factored x out of x^2+3x

OpenStudy (fibonaccichick666):

ok so now, what do you have to multiply (5x+20)/(x^2+3x) by to get a x+3 on the bottom, or what do you have to multiply (x-2)/(x+3) by to get a x^2+3x on the bottom?

OpenStudy (anonymous):

x?

OpenStudy (anonymous):

or 1/x

OpenStudy (fibonaccichick666):

ok I presonally prefer to make the common denominator x^2+3x here because who wants to deal with another fraction right?

OpenStudy (fibonaccichick666):

but so, you have to multiply by 1, you can't just multiply by x, so what do you have to multiply by?

OpenStudy (anonymous):

1/x? I'm confused

OpenStudy (fibonaccichick666):

ok, if you had 2/3+3/4 what would you do?

OpenStudy (anonymous):

multiply by 12

OpenStudy (fibonaccichick666):

well, you wouldn't multiply by 12

OpenStudy (fibonaccichick666):

you would have to get a common denominator of twelve, yes,

OpenStudy (fibonaccichick666):

but how would you add those, can you show me your work for that?

OpenStudy (anonymous):

Sorry I had to step away for a minute 2/3 + 3/4 1/4(2/3) + 1/3(3/4) 2/12 + 3/12 = 5/12

OpenStudy (fibonaccichick666):

it's ok, so there is an error in your work there, can you read this for me then tel me what it is? http://www.purplemath.com/modules/fraction4.htm

OpenStudy (anonymous):

4/4(2/3) + 3/3(3/4)

OpenStudy (anonymous):

8/12 + 9/12 = 17/12

OpenStudy (fibonaccichick666):

right so now, what does 4/4 simplify to?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

so do I multiply my problem by x/x?

OpenStudy (fibonaccichick666):

yes!

OpenStudy (anonymous):

progress! lol ok I'll work on that for a minute

OpenStudy (fibonaccichick666):

ok, just remember though it is only one fraction being multipled

OpenStudy (anonymous):

wait I messed up, I'll try again (5x+20)/(x^2+3x) - (x^2-2x)/(x^2+3x) = 6/x

OpenStudy (fibonaccichick666):

that's right

OpenStudy (fibonaccichick666):

ok so now, make the left one big old fraction

OpenStudy (anonymous):

(-x^2+3x+20)/(x^2+3x) = 6/x

OpenStudy (anonymous):

I messed up again :( (-x^2+7x+20)/(x^2+3x) = 6/x ??

OpenStudy (fibonaccichick666):

messing up is ok, let me look

OpenStudy (fibonaccichick666):

your second is correct :)

OpenStudy (fibonaccichick666):

I just get the feeling, are you a nontraditional student?

OpenStudy (anonymous):

yay!

OpenStudy (anonymous):

I do school online if that's what you mean

OpenStudy (fibonaccichick666):

oh ok, I thought you may be older returning to college type

OpenStudy (anonymous):

Oh no I'm only 17 haha

OpenStudy (fibonaccichick666):

lol ok, so anyways after here. you can get rid of the fraction on the right

OpenStudy (anonymous):

by multiplying with x/x again?

OpenStudy (anonymous):

or does it need the same denominator as the left side?

OpenStudy (fibonaccichick666):

not quite you just need to multiply by x to cancel out the x on the right

OpenStudy (anonymous):

(-x^2+7x+20)/(x^2+3x) = 6

OpenStudy (anonymous):

don't I still need to get rid of the fraction on the left? do I multiply everything by the denominator?

OpenStudy (fibonaccichick666):

not yet, first you have to make sure you multiply both sides by x

OpenStudy (anonymous):

i already multiplied the right side by x though

OpenStudy (fibonaccichick666):

sorry, had to go. but you are undoing the fraction 6/x, in order to do that you have to multiply both sides by x not just one

OpenStudy (anonymous):

that's fine! so then: x[(-x^2+7x+20)/(x^2+3x)]=(6/x)(x)

OpenStudy (fibonaccichick666):

yes

OpenStudy (anonymous):

(-x^3+7x^2+20x)/(x^2+3x) = 6

OpenStudy (anonymous):

now do I multiply both sides by x^2+3x?

OpenStudy (fibonaccichick666):

no, now you need to reduce the fraction

OpenStudy (anonymous):

I think I worked it out to the answer... x = 2 or x = -1

OpenStudy (anonymous):

and x = 0 was an extraneous solution

OpenStudy (fibonaccichick666):

x can't equal zero, what does your def of extraneous mena?

OpenStudy (fibonaccichick666):

mean?

OpenStudy (fibonaccichick666):

I have to go take a final but here is the rest, sorry. \[\frac{x(-x^2+7x+20)}{x(x+3)} = 6\]\[\frac{(-x^2+7x+20)}{(x+3)} = 6\]

OpenStudy (fibonaccichick666):

factor the top and some cancelling should happen